The molar concentration of a sugar solution in an open beaker has been determined to be 0.3M. Calculate the solute potential at 27 degrees. SHOW YOUR CALCULATIONS! Round your answer to the nearest hundredth. (Answers without shown work will not receive any credit)
1
Expert's answer
2020-02-03T10:18:52-0500

Solute potential can be calculated by the following formula:

Ψ = - M × i × R ×T

where M - the molar concentration of the solute, i - the van't Hoff factor, the ratio of amount of particles in solution to amount of formula units, R - the ideal gas constant, T - is the absolute temperature.

As M = 0.3 M, i = 1 (sugar molecules do not dissociate), R = 8.314 (L·kPa)/(mol·K), and T = 27 °C = 300.15 K:

Ψ (sugar) = 0.3 M × 8.314 (L·kPa)/(mol·K) × 300.15 K = 748.63 kPa


Answer: 748.63 kPa